OpenStudy (anonymous):

Bert is planning to open a savings account that earns 1.6% simple interest yearly. He wants to earn exactly $192 in interest after 2 years. How much money should he deposit? $3,072 $6,000 $60 $600

OpenStudy (anonymous):

@Chriss1234

OpenStudy (anonymous):

@cupcakerain

OpenStudy (anonymous):

i will give medal

OpenStudy (anonymous):

@help?!?!?!

OpenStudy (anonymous):

@help?!?!?!

OpenStudy (hotpinkgirl):

it is $6000

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

can you answer a few more?

OpenStudy (hotpinkgirl):

it says after two years so you do 1.6 * 2 = 3.2 $6000 + 3.2% = 6,192

OpenStudy (anonymous):

Dave rented a limousine for his wife's birthday. The hourly rate is $60. They used the limousine for 3 hours, plus Dave gave the driver a 10% tip. How much did he spend in total for the hourly charges plus tip? $162 $216 $246 $198

OpenStudy (hotpinkgirl):

your welcome

OpenStudy (anonymous):

@HotPinkGirl

OpenStudy (anonymous):

can you help with

OpenStudy (anonymous):

Dave rented a limousine for his wife's birthday. The hourly rate is $60. They used the limousine for 3 hours, plus Dave gave the driver a 10% tip. How much did he spend in total for the hourly charges plus tip? $162 $216 $246 $198

OpenStudy (anonymous):

180

OpenStudy (hotpinkgirl):

$60 * 3 hours = 180 + 10% = 198

OpenStudy (anonymous):

thx

OpenStudy (hotpinkgirl):

your welcome